OpenStudy (anonymous):

a right triangle has a leg length of 48 in and an hypotenuse of 80 in, what is the length of the other leg?

OpenStudy (anonymous):

use the pythagorean theorem: \(\large (hypotenuse)^2=(leg_1)^2+(leg_2)^2 \)

OpenStudy (anonymous):

lol I got 4096..

OpenStudy (anonymous):

I think it's 70 in

OpenStudy (anonymous):

or if you don't want to work with large numbers, notice that 48 = 16*3 and 80=16*5. they both have a common factor of 16 so just work with a leg of 3 and hypotenuse of 5... just be sure to multiply your answer by 16 at the end..

OpenStudy (anonymous):

is it 56 in?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

how did you get 56?

OpenStudy (anonymous):

while we're at it, how did you get 70?

OpenStudy (anonymous):

even if one of them is correct i need to see how you got it so i can say, "yes, it's correct"